1. Find x and y so that the ordered data set has a mean of 42 and a median of 35. 

17, 22, 26, 29,x,42,67,70,y


2. The exam grades of 7 students are given below. 70, 66, 72, 96, 46, 90, 50 

            Find 

  1. Mean 
  2. The standard deviation 
1
Expert's answer
2021-07-25T10:11:36-0400

Question 1:

It is important to note that this ordered data set has 9 elements; namely,

17, 22, 26, 29,x,42,67,70,y.

The median is the middle value when the data is arranged in ascending order. Therefore,

in this data the median is given as 35. x in the data represent the median since the number

of elements in this ordered data is odd. Therefore x=35

To get y, which is the last element, take mean multiply by 9 and then subtract the total elements given

y = (42 x 9) - (17 + 22+ 26+ 29+35 +42+67+70)

y = 70.

Question 2:

Mean = sum of the element / total number of elements

= (70+ 66+ 72+ 96+ 46+ 90+50 ) / 7

mean = 70

Standard Deviation = square root of variance

= σ2 = Σ(xi - μ)2 /N= (70 - 70)2 + ... + (50 - 70)27= 20727= 296σ = √296= 17.204650534085